The depth of a pool varies directly with how many gallons of water it will hold. If a 6 foot deep pool holds 10,000 gallons of water, a pool that holds 12,500 gallons of water should be __________inches deep, according to this theory.

Answer by checkley71(8403)

720/10000=X/12500
10000X=72*12500
10000X=9000000
X=9000000/10000
X=90 INCHES DEEP.